# version 1.0

### v6

@@ 3.2.1 Markov Chains
A Markov Chain (or Markov Process) is a model of a memoryless stochastic process as a weighted directed graph. This is the first layer in abstraction before allowing for actions, as it only describes a process without exogenous inputs.
It consists of a set of states and a transition matrix that describes the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$. Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as follows:
@@
Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chances of weather for the next day.
Denote the distribution over states at time $t$ by $\mathbf{p}_{t}$. Then it can be seen that
@@
where $\mathbf{p}_{0}$ is the initial state distribution represented as a vector.
A central object in the study of Markov Chains is stationary distributions.
@@
A stationary distribution of a Markov Process is a limiting probability distribution
$\boldsymbol{\pi}=\lim_{t\to\infty}\mathbf{p}_{0}P^{t}.$
A stationary distribution satisfies
@@
A stationary distribution does not always exist nor is it always unique. Existence, uniqueness, and convergence to stationarity are the subject of Ergodic theory, a rich and well-studied branch of mathematics. For our discussion, we are interested in sufficient conditions for stationary distributions to exist. They are:
@@
    Irreducibility. A Markov Chain is irreducible if and only if $\forall i,j\;\exists t\;s.t.\;P_{ij}^{t}>0$. In other words, every state is reachable from every other state.
@@
    Aperiodicity. A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where d(s_i) = gcd{t ∈N_+ | (P^t)_ii > 0 } , is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
The ergodic theorem states that every irreducible and aperiodic Markov chain has a unique stationary distribution. Its proof is a beautiful application of spectral graph theory. Exercise 3.6 asks the reader for a guided proof of the ergodic theorem.
@@
Figure 3.6: In this very simple Markov Chain, the state alternates every time step, so there is no unique stationary distribution. This Markov Chain has periodicity of two.
@@
Figure 3.7: Adding a self-loop to every node makes the periodicity to be one, and this chain is also irreducible. The stationary distribution exists and is unique, it is equal to $\boldsymbol{\pi}=(\frac{1}{2},\frac{1}{2})$.
Reversible Markov Chains
An interesting special case of Markov chains is the reversible Markov chain.
@@
A Markov Chain $P$ is said to be reversible with respect to distribution $\pi$ if and only if the following hold for every two states $i,j$:
@@
A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain is reversible.
In exercise 3.6, we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
@@ 3.2.2 Markov Reward Processes
Markov Reward Processes (MRPs) add the concept of reward to Markov Chains. In each state, the agent receives a reward, which, for simplicity, we assume to be in the range $[0,1]$. In terms of algorithmic consequences, no generality is lost due to this assumption, and this is left as an exercise.
The cumulative reward is the sum of discounted rewards over time, so that the immediate rewards are worth more than those received in the future. This is consistent with the economics literature, where a fixed amount of money is worth more today than in the future, because of positive interest rates.
@@ A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
  • $S$ is a set of states that can be finite or infinite.
  • $P\in{\mathbb{R}}^{S\times S}$ is the transition matrix.
  • $R\in[0,1]^{S\times S}$ is the reward function, where $R(s,s^{\prime})$ is the reward for moving from state $s$ to $s^{\prime}$.
  • $\gamma\in[0,1]$ is the discount factor.
We henceforth consider instantiations of Markov Reward Processes, and use the following notation:
@@
    We denote by $S_{t}\in S$ the random state of the process at time $t$.
@@
    $R_{t}$ = the random reward received at time $t$, starting from state $S_{t}$ and moving to $S_{t+1}$
@@
    $\hat{R}\in{\mathbb{R}}^{S}$ = the vector that for each state, denotes the expected reward in the next iteration as a function of the current state, ^R(s) = E[ R_t | S_t = s] .
@@
    $G_{t}=$ the random variable denoting the sum of the discounted rewards starting from time $t$ onwards. G_t = ∑_i=0^∞ R_t+i γ^i .
When the discount factor is strictly less than one, $G_{t}$ is well defined. In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that there is a stationary distribution for the underlying Markov chain, denoted by $\boldsymbol{\pi}$.
$G:=\lim_{t\to\infty}\frac{1}{t}\sum_{i=1}^{t}R_{i}=\sum_{s\in S}\boldsymbol{\pi}(s)\hat{R}(s)$
One of the most important definitions in reinforcement learning is the value function:
@@
The value function maps states to their expected infinite-horizon reward,
$\mathbf{v}(s)=\mathbb{E}[G_{t}|S_{t}=s].$
The value function says how valuable a state is as measured by the sum of its future discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
$\mathbf{v}(s)=\hat{R}(s)+\gamma\sum_{s^{\prime}}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
To see that this equation holds we can write,
$\mathbf{v}(s)$ $=\mathbb{E}[G_{t}|S_{t}=s]$
$=\mathbb{E}[\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}|S_{t}=s]$
$=\mathbb{E}[R_{t}+\gamma\sum_{i=0}^{\infty}\gamma^{i}R_{t+1+i}|S_{t}=s]$
$=\mathbb{E}[R_{t}+\gamma G_{t+1}|S_{t}=s]$
$=\mathbb{E}[R_{t}+\gamma\mathbf{v}(S_{t+1})|S_{t}=s]$
$=\hat{R}(s)+\gamma\sum_{s^{\prime}}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
In matrix form, we can write the Bellman equation as $\mathbf{v}=\mathbf{\hat{R}}+\gamma P\mathbf{v}$, which simplifies to $\mathbf{v}=(I-\gamma P)^{-1}\mathbf{\hat{R}}$. Given $P$ and $\mathbf{\hat{R}}$, this is a linear system of equations, and thus the value function can be solved using Gaussian elimination in cubic time in the number of states.
@@ 3.2.3 Markov Decision Processes
The final layer of abstraction is to allow actions or controls in the language of control theory. This gives the full Markov Decision Process model as defined below.
@@ A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
    $P$ is a transition matrix: P_ss’^a = Pr[ S_t+1 = s | S_t = s, A_t = a].
@@
    $R$ is a reward function that gives each transition and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$. We henceforth assume w.l.o.g. that $R\in[0,1]^{S\times A}$, where we define R_sa = _s’ S P_s,s’,a R_s,s’,a .
@@
The central solution concept of study in the theory of reinforcement learning is called a policy. A policy is a mapping from state to distribution over actions, that is,
$\pi:S\to\Delta(A)\ ,\ \pi(a|s)=\Pr[A_{t}=a|S_{t}=s]$
The goal of reinforcement learning is to determine a sequence of actions to maximize long-term reward. The Markovian structure of the problem suggests that this optimal value is realized by a stationary policy, i.e. there exists a fixed map from states to actions that maximizes the expected rewards.
Thus, the goal of reinforcement learning is to find a policy which maximizes the expected reward starting from a certain starting state, i.e.
$\max_{\pi\in S\to\Delta(A)}\mathop{\mbox{\bf E}}\left\{\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|a_{t}\sim\pi(s_{t})\ ,\ S_{1}=s_{1}\right\},$
where $R_{t}$ is a random variable that denotes the reward at time $t$ that depends on the randomness in the policy as well as the transitions.
Similarly to the value function of a Markov Reward Process, we can define the value function for a Markov Decision Process as follows.
The value function for a state $s$ given a policy $\pi$ is defined to be the expected reward from this state onwards where the actions are chosen according to $\pi$ at each state, i.e.
$\mathbf{v}_{\pi}(s)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s\right].$
Similarly, the $Q$ function is defined over state-action pairs given a policy $\pi$ to be the expected cumulative reward gained when starting from some state $s$ and action $a$ and thereafter choosing actions according to $\pi$ at each future state, i.e.
$Q_{\pi}(s,a)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s,A_{1}=a\right].$
A fundamental theorem of reinforcement learning, which is constructively proven in Section 3.4, is as follows.
@@
There exists a deterministic optimal policy $\pi^{*}$, which maps every state to a single action. All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for all states and actions.
The fact that there exists an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
@@ 3.3 The Bellman Equation
A fundamental property of Markov Decision Processes is that they satisfy the Bellman equation for dynamic programming. The importance of this equation is paramount in algorithm design: we describe two efficient methods for reinforcement learning based on this equation later in this chapter. In later chapters, we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
The Bellman equation for any policy $\pi$ relates the value function of that policy to itself through one step of application for that policy. Mathematically, this amounts to the following:
$\mathbf{v}_{\pi}(s)=\sum_{a\in A}\pi(a|s)\left[R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\mathbf{v}_{\pi}(s^{\prime})\right].$
It is sometimes convenient to use matrix-vector notation, where we can write:
@@
    $R^{\pi}$ is the expected reward vector induced when playing the policy $\pi$, that is, ${R^{\pi}}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(s)}[R_{sa}]$.
@@
    $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e. P^π_ss’ = ∑_a ∈A P^a_ss’ ⋅π(a|s).
The Bellman equation gains prominence when applied to the optimal policy of a Markov Decision Process. Then it is called the Bellman optimality equation, and states that for every state and action, we have
$\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}$
Equivalently, this equation can be written in terms of the $Q$ function as
$\mathbf{v}^{*}(s)=\max_{a}\left\{Q^{*}(s,a)\right\},$
where $Q^{*}$ is the $Q$ function of the optimal policy, i.e.,
$Q^{*}(s,a)=R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\max_{a^{\prime}}\left\{Q^{*}(s^{\prime},a^{\prime})\right\}.$
In subsequent chapters, we will also consider cases where the return of interest is either the reward accumulated in a finite number of time steps, that is, $G=\sum_{t=1}^{T}R_{t}$ in the finite-time setting, or the infinite-time average reward case, i.e. $G=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}R_{t}$.
In the finite-time setting, both optimal policies and the optimal value function are non-stationary – they depend on the time step in addition to the current state. In this setting, the corresponding Bellman optimality equation can be written as
$\mathbf{v}_{t}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}_{t+1}(s^{\prime})\right\}$
starting with $\mathbf{v}_{T}(s)=0$ for all states.
In the average-reward case, the Bellman optimality equation offers the following characterization: $\mathbf{v}^{*}:S\to{\mathbb{R}}$ is the optimal value function if there exists some $\lambda\in{\mathbb{R}}$ such that
$\lambda+\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
Here the scalar $\lambda$ corresponds to the optimal average long-term reward, and $\mathbf{v}^{*}(s)$ reflects the transient advantage of being at state $s$.
3.3.1 Reinforcement Learning as Linear Programming
One of the immediate conclusions from the Bellman optimality equation is an efficient algorithm for reinforcement learning in Markov Decision Processes via linear programming.
Consider the problem of computing the value function corresponding to an optimal policy. The optimal value function gives the optimal policy, and deriving this implication explicitly is left as an exercise.
According to the Bellman optimality equation, the optimal value function satisfies:
$\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\},$
or in vector form,
$\mathbf{v}^{*}=\max_{\mathbf{a}\in A^{|S|}}\left\{\mathbf{R^{\mathbf{a}}}+\gamma P^{\mathbf{a}}\mathbf{v}^{*}\right\},$
where the maximization over actions takes place independently for each state. Equivalently, we can write the following mathematical program:
$\min_{\mathbf{v}^{*}}\mathbf{1}^{\top}\mathbf{v}^{*}$
subject to $\mathbf{v}^{*}(s)\geq R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})$ $\forall(s,a)\in S\times A.$
Notice that this is a linear program! The number of variables is $|S|$, and the number of constraints is $|S|\times|A|$, which can be solved in polynomial time in the number of states and actions. The feasibility of this linear program is not immediate but will be established in the next section.

### v8

@@ 3.2.1 Markov Chains
A Markov chain is a stochastic process that evolves without exogenous actions. It consists of a finite set of states together with a transition matrix that specifies the probability of moving from one state to another.
Formally, let $S$ be a finite set of states. A Markov chain on $S$ is specified by a row-stochastic matrix $P\in[0,1]^{S\times S}$, where the entry $P_{ij}$ is the probability of moving from state $i$ to state $j$ in one step. Figure 3.5 describes a Markov chain corresponding to a simple weather process with transition matrix
@@
Figure 3.5: A graphical description of a Markov chain. Each state is the weather on a given day, and the transition probabilities describe the distribution of the next day’s weather.
Denote the distribution over states at time $t$ by $\mathbf{p}_{t}$. Then
@@
where $\mathbf{p}_{0}$ is the initial state distribution.
A central object in the study of Markov chains is a stationary distribution.
@@
A distribution $\boldsymbol{\pi}$} over the state space $S$ is a stationary distribution of a Markov chain with transition matrix $P$ if and only if
@@
For finite Markov chains, a stationary distribution always exists. It need not be unique, however, and convergence from an arbitrary initial distribution need not occur without further assumptions.
Two standard structural properties are the following.
@@
    Irreducibility. A Markov chain is irreducible if and only if
    $\forall i,j\in S\;\exists t\geq 0\text{ such that }(P^{t})_{ij}>0.$
    In words, every state is reachable from every other state.
@@
    Aperiodicity. A Markov chain is aperiodic if and only if every state has period one. The period of a state $s_{i}$ is defined by
    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \middle|\ (P^{t})_{ii}>0\right\}.$
The following finite-state ergodic theorem summarizes the role of these properties.
Theorem 3.2 (Ergodic theorem for finite Markov chains).
Let $P$ be a finite Markov chain.
  1. 1.
    $P$ has at least one stationary distribution.
  2. 2.
    If $P$ is irreducible, then the stationary distribution is unique.
  3. 3.
    If $P$ is irreducible and aperiodic, then for every initial distribution $\mathbf{p}_{0}$,
    $\lim_{t\to\infty}\mathbf{p}_{0}P^{t}={\boldsymbol{\pi}},$
    where $\boldsymbol{\pi}$} is the unique stationary distribution.
Its proof is a beautiful application of spectral graph theory. Exercise 3.6 guides the reader through one standard proof of the convergence statement.
@@
Figure 3.6: This two-state Markov chain is irreducible but periodic with period $2$. It has a unique stationary distribution, namely ${\boldsymbol{\pi}}=(\frac{1}{2},\frac{1}{2})$, but the distribution of the chain does not converge to $\boldsymbol{\pi}$} from every initial state because the chain alternates deterministically between the two states.
@@
Figure 3.7: Adding a self-loop to every node makes the period equal to one. This chain is irreducible and aperiodic, so its stationary distribution exists, is unique, and equals ${\boldsymbol{\pi}}=(\frac{1}{2},\frac{1}{2})$.
An important special case is the reversible Markov chain.
@@
A Markov chain $P$ is said to be reversible with respect to a distribution $\boldsymbol{\pi}$} if and only if, for every two states $i,j$,
@@
A Markov chain is said to be reversible if there exists some distribution with respect to which it is reversible.
Exercise 3.6 asks the reader to prove that any distribution satisfying the reversibility equations is automatically stationary.
@@ 3.2.2 Markov Reward Processes
Markov reward processes add rewards to Markov chains. At each time step, the process transitions between states and emits a reward. For simplicity, we assume rewards lie in the range $[0,1]$. This normalization loses no essential generality and keeps the formulas uncluttered.
@@ A Markov reward process is a tuple $(S,P,R,\gamma)$, where:
  • $S$ is a finite set of states.
  • $P\in[0,1]^{S\times S}$ is a row-stochastic transition matrix.
  • $R\in[0,1]^{S\times S}$ is a reward function, where $R(s,s^{\prime})$ is the reward obtained when moving from state $s$ to state $s^{\prime}$.
  • $\gamma\in[0,1)$ is the discount factor.
We henceforth use the following notation.
@@
    $S_{t}\in S$ denotes the random state of the process at time $t$.
@@
    $R_{t}$ denotes the random reward obtained when transitioning from $S_{t}$ to $S_{t+1}$.
@@
    $\hat{R}\in{\mathbb{R}}^{S}$ denotes the vector of expected one-step rewards:
    $\hat{R}(s)=\mathop{\mbox{\bf E}}[R_{t}\mid S_{t}=s].$
@@
    $G_{t}$ denotes the discounted return from time $t$ onward:
    $G_{t}=\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}.$
Since $\gamma<1$ and rewards are bounded, the random variable $G_{t}$ is well defined.
A natural question is: what quantity should summarize the future behavior of a Markov reward process? For an ordinary Markov chain, one often studies long-run state frequencies via stationary distributions. Once rewards and discounting are introduced, however, the central quantity becomes the expected discounted return from a given starting state. This dependence on the starting state is unavoidable: for example, when $\gamma=0$, the return reduces to the immediate reward, which can vary from state to state.
This motivates the central object in the theory of Markov reward processes: the value function. Because the process is time-homogeneous, the conditional expectation $\mathop{\mbox{\bf E}}[G_{t}\mid S_{t}=s]$ depends only on the state $s$, and not on the particular time index $t$.
@@
The value function of a Markov reward process maps states to expected discounted return:
$\mathbf{v}(s)=\mathbb{E}[G_{t}\mid S_{t}=s].$
The value function measures how desirable each state is in terms of expected future reward. It satisfies the Bellman equation for a Markov reward process:
$\mathbf{v}(s)=\hat{R}(s)+\gamma\sum_{s^{\prime}\in S}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
To verify this identity, write
$\mathbf{v}(s)$ $=\mathbb{E}[G_{t}\mid S_{t}=s]$
$=\mathbb{E}\left[\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}\;\middle|\;S_{t}=s\right]$
$=\mathbb{E}\left[R_{t}+\gamma\sum_{i=0}^{\infty}\gamma^{i}R_{t+1+i}\;\middle|\;S_{t}=s\right]$
$=\mathbb{E}[R_{t}+\gamma G_{t+1}\mid S_{t}=s]$
$=\mathbb{E}[R_{t}+\gamma\mathbf{v}(S_{t+1})\mid S_{t}=s]$
$=\hat{R}(s)+\gamma\sum_{s^{\prime}\in S}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
In matrix form, the Bellman equation is
$\mathbf{v}=\hat{R}+\gamma P\mathbf{v}.$
Equivalently, $\mathbf{v}$ is the unique solution to the linear system
$(I-\gamma P)\mathbf{v}=\hat{R}$.
Since $\gamma<1$ and $P$ is a stochastic matrix, the matrix $I-\gamma P$ is invertible, and this system has a unique solution.
Thus, for a known Markov reward process, the value function is the solution to a linear system and can be computed in polynomial time in the number of states.
@@ 3.2.3 Markov Decision Processes
The final layer of abstraction is to allow actions, or controls in the language of control theory. This gives the full Markov decision process model.
@@ A Markov decision process is a tuple $(S,P,R,A,\gamma)$ such that:
    $P$ is a transition kernel, written as
    $P^{a}_{ss^{\prime}}=\Pr[S_{t+1}=s^{\prime}\mid S_{t}=s,A_{t}=a].$
    For each action $a\in A$, the matrix $P^{a}$ is row-stochastic.
@@
    $R$ is a reward function on state–action–next-state triples, where $R_{ss^{\prime}a}$ is the reward for moving from $s$ to $s^{\prime}$ using action $a$. We also write
    $R_{sa}=\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}R_{ss^{\prime}a}$
    for the corresponding expected immediate reward. As above, we assume rewards lie in $[0,1]$.
@@
The central solution concept is a policy.
Definition 3.7.
A policy is a rule for choosing actions as the process evolves. A policy may in principle depend on the entire history. A stationary policy is a policy that depends only on the current state, and is therefore a mapping
$\pi:S\to\Delta(A),\qquad\pi(a\mid s)=\Pr[A_{t}=a\mid S_{t}=s].$
If each $\pi(\cdot\mid s)$ is a point mass, the policy is called deterministic.
In this chapter we consider infinite-horizon discounted return. Starting from a state $s$, the value of a policy $\pi$ is
$\mathbf{v}_{\pi}(s)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t}\;\middle|\;S_{0}=s\right].$
The optimal value function is defined pointwise by
$\mathbf{v}^{*}(s)=\sup_{\pi}\mathbf{v}_{\pi}(s).$
A fundamental theorem of discounted MDPs is the following.
@@
For every finite discounted Markov decision process, there exists a deterministic stationary optimal policy $\pi^{*}$. Moreover, every optimal policy has the same value function $\mathbf{v}^{*}$.
The existence of an optimal stationary policy is a consequence of the Markovian structure of the problem and the Bellman optimality equation, which we now discuss.
@@ 3.3 The Bellman Equation
A fundamental property of Markov decision processes is that they satisfy a recursive optimality principle. This principle is expressed by the Bellman equation and is the basis for dynamic programming.
For a fixed stationary policy $\pi$, the Bellman equation takes the form
$\mathbf{v}_{\pi}(s)=\sum_{a\in A}\pi(a\mid s)\left[R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\mathbf{v}_{\pi}(s^{\prime})\right].$
In matrix–vector notation this can be written as
@@
    $R^{\pi}$ is the expected reward vector induced by $\pi$, defined by
    $R^{\pi}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(\cdot\mid s)}[R_{sa}],$
@@
    $P^{\pi}$ is the transition matrix induced by $\pi$, defined by
    $P^{\pi}_{ss^{\prime}}=\sum_{a\in A}P^{a}_{ss^{\prime}}\,\pi(a\mid s).$
Since $\gamma<1$, the matrix $I-\gamma P^{\pi}$ is invertible, and therefore the Bellman equation for a fixed policy has a unique solution.
The Bellman equation becomes especially powerful when applied to the optimal value function.
$\mathbf{v}^{*}(s)=\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}$
This is the Bellman optimality equation. It says that the optimal value of a state is obtained by choosing the action with the largest one-step reward plus discounted continuation value.
The proof is a one-step decomposition. Fix a state $s$. For any policy, the value obtained from $s$ is bounded above by the best possible choice of first action plus the optimal continuation value thereafter, so
$\mathbf{v}^{*}(s)\leq\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
Conversely, choose an action $a^{*}$ achieving the maximum on the right-hand side. For each successor state $s^{\prime}$, choose a policy that is $\varepsilon$-optimal from $s^{\prime}$. By first taking action $a^{*}$ and then following the chosen continuation policy, one achieves a value at least
$R_{sa^{*}}+\gamma\sum_{s^{\prime}\in S}P^{a^{*}}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime}),$
and the reverse inequality follows.
The Bellman optimality equation immediately implies the existence of a deterministic optimal policy. Indeed, define a policy $\pi^{*}$ by choosing at each state any maximizing action,
$\pi^{*}(s)\in\arg\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
Then $\mathbf{v}^{*}$ satisfies
$\mathbf{v}^{*}(s)=R_{s\pi^{*}(s)}+\gamma\sum_{s^{\prime}\in S}P^{\pi^{*}(s)}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\qquad\text{for all }s\in S,$
which is exactly the Bellman equation for the fixed policy $\pi^{*}$. By uniqueness of the solution of the Bellman equation for a fixed policy, we obtain $\mathbf{v}_{\pi^{*}}=\mathbf{v}^{*}$, proving that $\pi^{*}$ is optimal. This proves Theorem 3.8.
3.3.1 Linear Programming Formulation
The Bellman optimality equation characterizes $v^{*}$ as a fixed point. An equivalent characterization views $v^{*}$ as the smallest function satisfying the Bellman inequalities.
Consider the following linear program:
$\min_{\mathbf{v}\in\mathbb{R}^{|S|}}\quad$ $\sum_{s\in S}\mathbf{v}(s)$
s.t. $\mathbf{v}(s)\geq r(s,a)+\gamma\sum_{s^{\prime}\in S}P(s^{\prime}\mid s,a)\,\mathbf{v}(s^{\prime})\quad\forall s\in S,\ a\in A.$
This program has $|S|$ variables and |S||A||S|\,|A| constraints, so it can be solved in polynomial time in the numbers of states and actions. The next theorem shows that its unique optimum is precisely the optimal value function.
Theorem 3.9.
The unique optimal solution of the above linear program is the optimal value function $\mathbf{v}^{*}$.
Proof.
First, $\mathbf{v}^{*}$ is feasible, since it satisfies the Bellman optimality equation and therefore all Bellman inequalities.
Now let $\mathbf{v}$ be any feasible solution. Then for all $s$,
$\mathbf{v}(s)\geq\max_{a}\left\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}\mid s,a)\,\mathbf{v}(s^{\prime})\right\}=(T\mathbf{v})(s),$
where $T$ is the Bellman optimality operator. Thus $\mathbf{v}\geq T\mathbf{v}$. Iterating,
$\mathbf{v}\geq T\mathbf{v}\geq T^{2}\mathbf{v}\geq\cdots.$
Since $T^{k}\mathbf{v}\to\mathbf{v}^{*}$ as $k\to\infty$, it follows that $\mathbf{v}\geq\mathbf{v}^{*}$ componentwise.
Therefore $\mathbf{v}^{*}$ is the smallest feasible vector, and hence minimizes any strictly increasing objective such as $\sum_{s}\mathbf{v}(s)$. Uniqueness follows. ∎